OpenStudy (anonymous):

1/x^-2 for x=7 a. -49 b. 49

OpenStudy (anonymous):

\[\large \frac{1}{x^{-a}} = x^a\] Use this concept..

OpenStudy (anonymous):

so is it 49?

OpenStudy (anonymous):

Yes absolutely right.. Well done..

OpenStudy (anonymous):

thanks again!! ;)

OpenStudy (anonymous):

Welcome dear..

OpenStudy (anonymous):

And remember this concept also: \[\large \frac{1}{x^a} = x^{-a}\]